3x+y=0
-3x+y=-6
please help i'm very confusted as how to do this type of problem
thanks

Answer by narayaba(40)

3x + y = 0 equation 1
-3x + y = -6 equation 2
from equation 1 we get y = -3x
plugging in for -3x in equation 2
y + y =-6
2y =-6
y = -3
-3x = y
-3x =-3
x =1
x= 1 and y =-3